OpenStudy (anonymous):

Find an equation for the nth term of a geometric sequence where the second and fifth terms are -2 and 16, respectively. a. an = 1 • (-2)n - 1 b. an = 1 • 2n c. an = 1 • (-2)n + 1 d. an = 1 • 2n - 1

hartnn (hartnn):

know the formula for n'th term of geometric sequence ?

OpenStudy (anonymous):

Is it: a(n) = a*r^n-1 ?

hartnn (hartnn):

yes! \(\Large a_n = a_1 r^{n-1}\)

OpenStudy (anonymous):

Okay. So where do I go from there?

hartnn (hartnn):

so for 2nd term, put n =2 and a2 = -2 what do u get ?

OpenStudy (anonymous):

\[a _{n} = -2r ^{2-1}\] Is that what you mean?

OpenStudy (ilovecake):

hi hi hi

hartnn (hartnn):

but we don't know a_1 -2 is actually a_2 , not a_1

hartnn (hartnn):

\(\large a_2 = a_1 r^{2-1} \\ -2 =a_1 r \) got this ?

OpenStudy (anonymous):

So would I be using a different formula then?

OpenStudy (anonymous):

Oh okay.

OpenStudy (anonymous):

Yeah, I got it. One sec...

OpenStudy (anonymous):

Okay so now that we have that formula, what do I do? How do I know what r is?

hartnn (hartnn):

we also have 5th term! put n=5 and a5 = 16 in the same formula

OpenStudy (anonymous):

\[16 = a1^{r4}\]

OpenStudy (anonymous):

Is that wrong or completely wrong?

hartnn (hartnn):

\(\Large 16 = a_1 r^4\) thats correct! now we need to find 'r' know how to find r from these 2 equations ??? hint : divide

OpenStudy (anonymous):

16/4 = 4 r = 4

OpenStudy (anonymous):

Is that right or did I skip steps?

hartnn (hartnn):

-2 = a1 r 16 = a1 r^4 dividing -2/16 = r/r^4 got this ?

hartnn (hartnn):

note we divided , so that a1 gets cancelled and we could get r

OpenStudy (anonymous):

so r = 8?

OpenStudy (anonymous):

or is it r = -2

hartnn (hartnn):

r^3 = -8 so, yeah, r= -2 :)

OpenStudy (anonymous):

Okay :) So what's next?

hartnn (hartnn):

-2 = a_1 r you know r find a1

OpenStudy (anonymous):

How do I apply all of that into this?: a. an = 1 • (-2)n - 1 b. an = 1 • 2n c. an = 1 • (-2)n + 1 d. an = 1 • 2n - 1

hartnn (hartnn):

coming to it :)

OpenStudy (anonymous):

Okay.

OpenStudy (anonymous):

-2 = a1 r and r = - 2 so does a1 = 1? or 0?

hartnn (hartnn):

-2 = a1 * (-2) a1 = -2/-2 = 1 :) now plug the values in general formula! \(\Large a_n = a_1 r^{n-1}\)

OpenStudy (anonymous):

\[a _{n} = 1(-2) a _{n} = -2\]

OpenStudy (anonymous):

an = -2 correct?

hartnn (hartnn):

keep an as an \(\Large a_n = 1 \times (-2)^{n-1}\) is than an option ?

OpenStudy (anonymous):

Oh okay. And yes, it is. Thanks so much for helping me!

hartnn (hartnn):

welcome ^_^

OpenStudy (ilovecake):

suppppsssssssssssssssssss